A bucket full of water and weighing 200N is supported by two men holding the handle at each side, each pulling sideways at an angle of 30° to the vertical. Find the force in each man's arm.

1
Expert's answer
2022-03-08T08:55:58-0500

At the equilibrium

"mg=2F\cos30^\circ"

The force in each man's arm

"F=\frac{mg}{2\cos30^\circ}=\frac{200\:\rm N}{2*\sqrt{3}/2}=115\: \rm N"